OpenStudy (anonymous):

P(A) = .6, P(B) = .3, P(B | A) = .5

OpenStudy (anonymous):

P(A and B) = A. .06 B. .30 C. .03 D. .18 E. .05

OpenStudy (r.hamza17):

Is this 18 or 0.18?

OpenStudy (anonymous):

0.18

OpenStudy (r.hamza17):

Is there any other data for this question?

OpenStudy (anonymous):

No, just that one at the top

OpenStudy (zarkon):

\[P(B|A)=\frac{P(A\text{ and }B)}{P(A)}\] then \[P(A\text{ and }B)=P(B|A)P(A)\]

OpenStudy (anonymous):

So at first, it would be, \[P(B|A) = \frac{ .9 }{ .6 }\] ? Or did I do that wrong?

OpenStudy (r.hamza17):

Here is a method to solve it; Probability that A occurs is P(A)=.6 Probability that B occurs is P(B)=.3 Probability that event A does not occur, P(A')=0.4 Probability that event B does not occur, P(B')=0.7 Probability that event A and event B both occur, P(A∩B)=0.18

OpenStudy (zarkon):

\[P(A\text{ and }B)=P(B|A)P(A)=.5\times.6\]

OpenStudy (anonymous):

Ah, I see. The solution would be .30?

OpenStudy (zarkon):

yes

OpenStudy (anonymous):

Thank you so much!